what would be the slope of (-16, -10) and (-20, -13)? Thank you so much, I should have this by now but for some reason it is not sticking.
Found 2 solutions by mananth, stanbon:
Answer by mananth(16949)

x1 y1 x2 y2
-16 -10 -20 -13
(-16, -10) and (-20, -13)
slope m =(y2-y1)/(x2-x1)
(-13 -(-10) )/(-20 -(-16))
(-3/-4)
m=0.75

Answer by stanbon(75887)

what would be the slope of (-16, -10) and (-20, -13)
------------------
slope + (-13--10)/(-20--16) = -3/-4 = 3/4
============================================
Cheers,
Stan H.
=========